If the points (3,2) and (x,-5) lie on a line whose slope is (-7/2), then x equals



--i graphed it to see if i could do it and got x=1, but i got the answer wrong, please help.

Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)
Answer by Alan3354(69443)

If the points (3,2) and (x,-5) lie on a line whose slope is (-7/2), then x equals
----------------
slope, m, = diffy/diffx
-7/2 = (7)/(3-x)
1/2 = -1/(3-x)
3-x = -2
x = 5